# BioNeRF: Biologically Plausible Neural Radiance Fields for View Synthesis

### v2

@@ Abstract
keywords:
View Synthesis, Neural Rendering, iologically Plausible Neural Models
@@ 1 Introduction
Neural Radiance Fields (NeRF) mildenhall2021nerf provide a memory-efficient way to address the problem of rendering new synthetic views based on a set of input images and their respective camera poses. The model implements a Multilayer Perceptron (MLP) network whose weights store a tridimensional scene representation, producing photorealistic quality outputs from new viewpoints while preserving complex features concerning material reflectance and geometry.
Roughly speaking, NeRF comprises a fully connected network that performs regression tasks, i.e., it maps a 5D coordinate sampled from a particular camera viewpoint location relative to the scene, $(x,y,z)$, and their corresponding 2D viewing directions $(\theta,\phi)$, to a view-dependent RGB color, comprising a set of three real-valued numbers, and a single volume density $\Delta$. Such 5D coordinates are transformed using a positional encoding approach to extract higher frequencies, consequently improving the quality of high-resolution representations. Further, the model computes the radiance emitted in such direction using classical volume rendering techniques kajiya1984ray , accumulating those colors and densities into a 2D image. Figure 1 illustrates the concept behind NeRF.
Figure 1: NeRF synthesizes images by sampling 5D coordinates (location and viewing direction) along camera rays (left) and feeding those locations into an MLP to produce color and volume density (right).
One of NeRF’s fundamental concepts regards encoding complex real-world 3D scene representation in the parameters of a neural network. Besides, the model restricts the volume density prediction as a function of the camera position, allowing the RGB color to be predicted as a function of both location and viewing direction. Such concepts resemble some more biologically plausible studies inspired by neuroscience discoveries and principles of pyramidal cells kording2001supervised , especially concerning the idea of context to steer the information flow adeel2020conscious ; kay1998contextually ; passos2023multimodal and integrated memory, responsible for providing additional context based on past experiences adeel2022context ; passos2023canonical .
@@
The remainder of this paper is described as follows. Section 2 presents the related works, while Section 3 provides a theoretical background regarding the NeRF concepts. Section 4 introduces the BioNeRF, while Sections 5 and 6 describe the methodology and present the experimental results and discussions, respectively. Finally, Section 7 states conclusions and future works.
@@ 2 Related Works
Before NeRF, Mildenhall et al. mildenhall2019local proposed the Local Light Field Fusion (LLFF), a method for synthesizing high-quality scene views requiring a significantly smaller number of instances than traditional methods at the time. Meanwhile, Lombardi et al. lombardi2019neural proposed Neural Volumes (NV), a neural volumetric representation method for modeling and rendering dynamic scenes. The technique improved the image resolution without increasing the resolution of the voxel grid. Besides, Sitzmann et al. sitzmann2019scene proposed the Scene Representation Networks (SRNs), which employed unsupervised learning and convolutional neural networks to learn rich representations of 3D scenes without explicit supervision. Further, Liu et al. liu2020neural proposed the Neural Sparse Voxel Fields (NSVF), which rendered dynamic, large-scale scenes using voxel representations and progressive training to obtain high-quality real-time renderings with a reduced time. Finally, in 2021, Mildenhall et al. mildenhall2021nerf proposed the Neural Radiance Fields. This method uses a fully connected neural network to represent scenes as neural radiance fields, synthesizing new views of complex scenes given the camera position and direction and predicting the volumetric densities and emitted color, combined to generate new views using traditional rendering techniques.
@@
In the same year, Wang et al. wang2021ibrnet proposed the Image-based Rendering Neural Radiance Field (IBRNet), a learning-based image rendering method whose architecture consists of a multilayer perceptron network and a ray transformer, capable of continuously predicting colors and spatial densities from multiple views and a per-scene fine-tuning procedure. In the meantime, Barron et al. barron2021mip introduced the Multiscale Representation for Anti-Aliasing Neural Radiance Fields (Mip-NeRF). This method extends NeRF to represent the scene at a continuously-valued scale, improving the representation and rendering of three-dimensional scenes and overcoming the limitations of aliasing and blurring.
@@
Chen et al. chen2022tensorf proposed the Tensorial Radiance Fields (TensoRF), a technique that models and reconstructs radiance fields using 4D tensor representation. The work employs the CANDECOMP/PARAFAC decomposition to factorize tensors into compact components and vector-matrix decompositions to relax the component’s constraints. Concurrently, Fridovich et al. fridovich2022plenoxels presented the Radiance fields without neural networks (Plenoxels), a system for photorealistic image synthesis that represents a scene as a sparse 3D grid with spherical harmonics. The work of Xu et al. xu2022point introduced the Point-Based Neural Radiance Fields (Point-NeRF), a method for reconstructing and rendering three-dimensional scenes that includes rebuilding a point directly from input images via network inference. Furthermore, Verbin et al. verbin2022ref focused on representing specular reflections in 3D scenes in the Structured View-Dependent Appearance for Neural Radiance Fields (Ref-NeRF). The model employs structured refinements to capture vision-dependent appearance, obtaining state-of-the-art results with photorealistic images generated from new points of view.
@@
More recently, Chen et al. chen2023local proposed the Local-to-Global Registration for Bundle-Adjusting Neural Radiance Fields (L2G-NeRF). This method combines local pixel and global frame alignment to achieve high-fidelity reconstruction and addresses large camera pose misalignments. Meanwhile, Kulhanek et al. kulhanek2023tetra presented the Tetra-NeRF, a method for representing neural radiance fields using tetrahedra that combines concepts from 3D geometry processing, triangle-based rendering, point cloud generation, Delaunay triangulation, barycentric interpolation, and modern neural radiance fields. Finally, Yao et al. yao2023spiking proposed the SpikingNeRF, which aligns the radiance rays with the temporal dimension of Spiking Neural Networks, producing an energy-efficient approach associated with biologically plausible inspired concepts.
3 Neural Radiance Fields
Mildenhall et al.mildenhall2021nerf proposed NeRF for view synthesis purposes, i.e., generating 3D scenes or objects given a set of images captured at different angles. The process consists of approximating a function $F_{\Theta}:(\mathbf{l},\mathbf{d})\rightarrow(\mathbf{c},\Delta)$ that receives the object’s 3D location as input, i.e., $\mathbf{l}=(x,y,z)$, and its 2D viewing direction $\mathbf{d}=(\theta,\phi)$. The function outputs the object’s emitted color $\mathbf{c}=(r,g,b)$ and its volumetric density $\Delta$. We can represent function $F_{\Theta}$ using an MLP in which one wants to optimize the weights $\Theta$.
The MLP network processes the input $\mathbf{l}$ through 8 fully connected layers using a Rectified Linear Unit (ReLU) activation function and 256 channels per layer, and the output consists of the $\Delta$ and a feature map with 256 dimensions. At this stage, $\Delta$ does not depend on the view direction, i.e., a translucent point remains translucent regardless of the direction in which it is viewed. Next, the feature vector is concatenated with $\mathbf{d}$, which represents the view direction and serves as input to a fully connected layer using ReLU and 128 channels, which outputs the color information $\mathbf{c}$.
Rays passing through the scene are processed using the principles of volume rendering, computed through differentiability, which can be optimized using the gradient descent method. The objective is to minimize the error between the observed image and the rendered view.
One can summarize NeRF working mechanism as follows:
  • 1.
    To cast camera rays through the scene to sample 3D points;
  • 2.
    To use the 3D points $\mathbf{l}$ and their corresponding 2D viewing directions $\mathbf{d}$ as input to an MLP to produce a set of output colors $\mathbf{c}$ and densities $\Delta$; and
  • 3.
    To employ volume rendering techniques to accumulate those colors and densities into a 2D image.
@@ 4 Biologically Plausible Neural Radiance Fields
This paper introduces BioNeRF, which combines features extracted from both the camera position and direction into a memory mechanism inspired by cortical circuits canonicalLaminar ; canonicalCortical . Further, such memory is employed as a context to leverage the information flow into deeper layers. Figure 2 illustrates the BioNeRF architecture, which spans four main modules: (i) Positional Feature Extraction, (ii) Cognitive Filtering, (iii) Memory Updating, and (iv) Contextual Inference.
@@
Cognitive Filtering. This step performs a series of operations from now on called filters that work on the embeddings coming from the previous step [^1] We call those operations filters, for they output a value in the interval $[0,1]$ for each feature coming from the embeddings $\bm{h}_{\Delta}$ and $\bm{h}_{c}$.. There are four filters this step derives: (i) density ($\bm{f}_{\Delta}$), color ($\bm{f}_{c}$), memory ($\bm{f}_{\psi}$), and modulation ($\bm{f}_{\mu}$), computed as follows:
$\bm{f}_{\Delta}=\sigma\left(\bm{h}_{\Delta}\right),$ (1)
$\bm{f}_{c}=\sigma\left({\bm{h}_{c}}\right),$ (2)
$\bm{f}_{\psi}=\sigma\left({\bm{W}_{\psi}\bm{[h}_{\Delta},\bm{h}_{c}]+\bm{b}_{\psi}}\right),$ (3)
@@
$\bm{f}_{\mu}=\sigma\left({\bm{W}_{\mu}[\bm{h}_{\Delta},\bm{h}_{c}]+\bm{b}_{\mu}}\right),$ (4)
where $\bm{W}_{\psi}$ and $\bm{W}_{\mu}$ correspond to the weight matrices for the memory and modulation filters, respectively. Additionally, $\bm{b_{\psi}}$ and $\bm{b_{\mu}}$ stand for their respective biases. Moreover, $[\bm{h}_{\Delta},\bm{h}_{c}]$ represents the concatenation of embeddings $\bm{h}_{\Delta}$ and $\bm{h}_{c}$, while $\sigma$ denotes a sigmoid function. The pre-modulation $\bm{\gamma}$ is computed as follows:
@@
Loss Function. Let $r:\Re\times\Re^{3}\rightarrow\Re^{3}$ be a volume rendering technique kajiya1984ray that computes the pixel color given the volume density and the color. The BioNeRF loss function is defined as follows:
@@
The error is then back-propagated to the model’s previous layers/steps to update its set of weights ($\bm{W}=\{\bm{W}_{\Delta},\bm{W}_{c},\bm{W}_{\psi}$,$\bm{W}_{\mu},\bm{W}_{\gamma},\bm{W}_{\Psi},\bm{W}^{\prime}_{\Delta},\bm{W}^{\prime}_{c}\}$) and biases $\bm{b}=\{\bm{b}_{\Delta},\bm{b}_{c},\bm{b}_{\psi}$,$\bm{b}_{\mu},\bm{b}_{\gamma},\bm{b}_{\Psi},\bm{b}^{\prime}_{\Delta},\bm{b}^{\prime}_{c}\}$.
@@ 5.1 Datasets
We conducted experiments over two well-known datasets concerning view synthesis, i.e., Realistic Synthetic ${}^{\circ}$ mildenhall2021nerf and LLFF mildenhall2019local :
5.1.1 Realistic Synthetic ${}^{\circ}$.
Also known as Blender or NeRF Synthetic, the Realistic Synthetic ${}^{\circ}$ comprises eight object scenes with intricate geometry and realistic non-Lambertian materials. Six of these objects are rendered from viewpoints tested on the upper hemisphere, while the remaining two come from viewpoints sampled on a complete sphere, all at a resolution of $800\times 800$ pixels. Figure 3 (top) displays an instance image from each scene that compose the dataset.
5.1.2 Local Light Field Fusion (Real).
This paper considers $8$ scenes from the LLFF Real dataset, which comprises $24$ scenes captured from handheld cellphones with $20-30$ images each. The authors used a COLMAP structure from motion schonberger2016structure implementation to compute the poses. Figure 3 (bottom) provides a sample image from the scenes used in this work concerning this dataset.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
chair drums ficus hotdog lego materials mic ship
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
fern flower fortress horns leaves orchids room trex
Figure 3: Scene sample images extracted from Realistic Synthetic ${}^{\circ}$ (top) and LLFF (bottom) datasets.
@@ 5.2 Experimental Setup
In the sequence, the memory is used as a context to the consecutive blocks, being concatenated with the camera 3D coordinates to feed the second neural model $M^{\prime}_{\Delta}$, which comprises two dense layers with $256$ neurons and an output layer with a single neuron to predict the volume density $\Delta$. Similarly, the memory is concatenated with the 2D viewing direction $(\theta,\phi)$ to feed the second model $M^{\prime}_{c}$, which is composed of a dense layer with $128$ neurons and an output layer with three neurons to predict the directional emitted color. The network parameters were selected empirically based on the methodology employed in mildenhall2021nerf , consisting of the Adam optimizer with a learning rate of $45e-$ 4 during $400k$ updates.
@@
The experiments were conducted on an Ubuntu 18 system running on a 2x Intel® Xeon Bronze 3104 processor with $62$GB of memory and an NVIDIA® Tesla T4 GPU. The code was implemented using Python and the PyTorch framework and is available on GitHub [^2] https://github.com/Leandropassosjr/BioNeRF..
@@ 6.1 General Evaluation
Table 1 provides the general quantitative results for view synthesis, comparing BioNeRF against state-of-the-art methods considering both synthetic and real image datasets. The analysis is conducted based on three image reconstruction metrics: PSNR and SSIM, whose objective is to be maximized, and LPIPS, whose lower values denote better results. Notice that bolded numbers represent the most accurate results.
Table 1: Quantitative results considering both synthetic and real image datasets.
Realistic Synthetic ${}^{\circ}$mildenhall2021nerf LLFF (Real)mildenhall2019local
Method PSNR normal-$\uparrow$ SSIM normal-$\uparrow$ LPIPS normal-$\downarrow$ PSNR normal-$\uparrow$ SSIM normal-$\uparrow$ LPIPS normal-$\downarrow$
SRN sitzmann2019scene (NeurIPS’19) $22.26$ $0.846$ $0.170$ $22.84$ $0.668$ $0.378$
NV lombardi2019neural (ACM ToG 19) $26.05$ $0.893$ $0.160$- -- --
LLFF mildenhall2019local (ACM ToG 2019) $24.88$ $0.911$ $0.114$ $24.13$ $0.798$ $0.212$
NSVF liu2020neural (NeurIPS’20) $31.75$ $0.964$ $0.047$- -- --
NeRF mildenhall2021nerf (ACM 21) $31.01$ $0.947$ $0.081$ $26.50$ $0.811$ $0.250$
IBRNet wang2021ibrnet (CVPR’21) $28.14$ $0.942$ $0.072$ $26.73$ $0.851$ $0.175$
Mip-NeRF barron2021mip (ICCV’21) $33.09$ $0.961$ $0.043$- -- --
TensoRF chen2022tensorf (ECCV’22) $33.14$ $0.963$ $0.027$ $26,73$ $0.839$ $0.124$
Plenoxels fridovich2022plenoxels (CVPR’22) $31.71$ $0.958$ $0.049$ $26,29$ $0.839$ $0.210$
Point-NeRF xu2022point (CVPR’22) $33.31$ $0.978$ $0.027$- -- --
Ref-NeRF verbin2022ref (CVPR’22) $\bm{33.99}$ $0.966$ $0.038$- -- --
L2G-NeRF chen2023local (CVPR’23) $28.62$ $0.930$ $0.070$ $24.54$ $0.750$ $0.200$
Tetra-NeRF kulhanek2023tetra (ICCV’23) $32.53$ $\bm{0.982}$ $0.041$- -- --
SpikingNeRF yao2023spiking (arXiv’23) $32.45$ $0.956$- -- -- --
@@
22footnotetext: Values marked with -- describe results not provided by the authors.
The results confirm the effectiveness of BioNeRF since it could obtain the best outcomes for all three metrics considering the LLFF dataset, which comprises $8$ scenes from a real environment. BioNeRF also obtained the lowest LPIPS value over the synthetic dataset, outperforming all state-of-the-art results in this context. Such results are highly positive since LPIPS correspond to human perceptual judgments far better than the commonly used metrics like PSNR and SSIM zhang2018unreasonable .
@@
Table 2: Quantitative results considering each scene from Realistic Synthetic ${}^{\circ}$ dataset.
PSNR normal-$\uparrow$

### v3

@@ 1 Introduction
Neural Radiance Fields (NeRF) [16] provide a memory-efficient way to address the problem of rendering new synthetic views based on a set of input images and their respective camera poses. The model implements a Multilayer Perceptron (MLP) network whose weights store a tridimensional scene representation, producing photorealistic quality outputs from new viewpoints while preserving complex features concerning material reflectance and geometry.
Roughly speaking, NeRF comprises a fully connected network that performs regression tasks, i.e., it maps a 5D coordinate sampled from a particular camera viewpoint location relative to the scene, $(x,y,z)$, and their corresponding 2D viewing directions $(\theta,\phi)$, to a view-dependent RGB color, comprising a set of three real-valued numbers, and a single volume density $\Delta$. Such 5D coordinates are transformed using a positional encoding approach to extract higher frequencies, consequently improving the quality of high-resolution representations. Further, the model computes the radiance emitted in such direction using classical volume rendering techniques [11], accumulating those colors and densities into a 2D image.
One of NeRF’s fundamental concepts regards encoding complex real-world 3D scene representation in the parameters of a neural network. Besides, the model restricts the volume density prediction as a function of the camera position, allowing the RGB color to be predicted as a function of both location and viewing direction. Such concepts resemble some more biologically plausible studies inspired by neuroscience discoveries and principles of pyramidal cells [13], especially concerning the idea of context to steer the information flow [1, 12, 17] and integrated memory, responsible for providing additional context based on past experiences [2, 18].
@@
The remainder of this paper is described as follows. Section 2 presents the related works. Further, Section 3 introduces the BioNeRF, while Sections 4 and 5 describe the methodology and present the experimental results and discussions, respectively. Finally, Section 6 states conclusions and future works.
@@ 2 Related Works
Before NeRF, Mildenhall et al. [15] proposed the Local Light Field Fusion (LLFF), a method for synthesizing high-quality scene views requiring a significantly smaller number of instances than traditional methods at the time. Besides, Sitzmann et al. [21] proposed the Scene Representation Networks (SRNs), which employed unsupervised learning and convolutional neural networks to learn rich representations of 3D scenes without explicit supervision.Finally, in 2021, Mildenhall et al. [16] proposed the Neural Radiance Fields. This method uses a fully connected neural network to represent scenes as neural radiance fields, synthesizing new views of complex scenes given the camera position and direction and predicting the volumetric densities and emitted color, combined to generate new views using traditional rendering techniques.
@@
In the same year, Wang et al. [23] proposed the Image-based Rendering Neural Radiance Field (IBRNet), a learning-based image rendering method whose architecture consists of a multilayer perceptron network and a ray transformer, capable of continuously predicting colors and spatial densities from multiple views and a per-scene fine-tuning procedure. In the meantime, Barron et al. [3] introduced the Multiscale Representation for Anti-Aliasing Neural Radiance Fields (Mip-NeRF). This method extends NeRF to represent the scene at a continuously-valued scale, improving the representation and rendering of three-dimensional scenes and overcoming the limitations of aliasing and blurring. Further, the same authors proposed Mip-NeRF 360 [4], an extension of Mip-NeRF that employs online distillation, non-linear scene parameterization, and distortion-based regularizer to tackle problems related to unbounded scenes present in the prior version.
@@
Chen et al. [6] proposed the Tensorial Radiance Fields (TensoRF), a technique that models and reconstructs radiance fields using 4D tensor representation. The work employs the CANDECOMP/PARAFAC decomposition to factorize tensors into compact components and vector-matrix decompositions to relax the component’s constraints. Concurrently, Fridovich et al. [8] presented the Radiance fields without neural networks (Plenoxels), a system for photorealistic image synthesis that represents a scene as a sparse 3D grid with spherical harmonics. The work of Xu et al. [24] introduced the Point-Based Neural Radiance Fields (Point-NeRF), a method for reconstructing and rendering three-dimensional scenes that includes rebuilding a point directly from input images via network inference. Furthermore, Verbin et al. [22] focused on representing specular reflections in 3D scenes in the Structured View-Dependent Appearance for Neural Radiance Fields (Ref-NeRF). The model employs structured refinements to capture vision-dependent appearance, obtaining state-of-the-art results with photorealistic images generated from new points of view.
@@
Chen et al. [7] proposed the Local-to-Global Registration for Bundle-Adjusting Neural Radiance Fields (L2G-NeRF). This method combines local pixel and global frame alignment to achieve high-fidelity reconstruction and addresses large camera pose misalignments. Meanwhile, Kulhanek et al. [14] presented the Tetra-NeRF, a method for representing neural radiance fields using tetrahedra that combines concepts from 3D geometry processing, triangle-based rendering, point cloud generation, Delaunay triangulation, barycentric interpolation, and modern neural radiance fields. Meanwhile, Yao et al. [25] proposed the SpikingNeRF, which aligns the radiance rays with the temporal dimension of Spiking Neural Networks, producing an energy-efficient approach associated with biologically plausible inspired concepts. Finally, Selvaratnam et al. [20] proposed the Localised-NeRF, which projects pixel points onto training images to obtain representation on HSV gradient space and further extract features to synthesize novel views. The method employs an attention-driven approach to maintain direction consistency and leverage image-based features to predict the diffuse and specular colors, exhibiting competitive performance with prior NeRF-based models.
@@ 3 Biologically Plausible Neural Radiance Fields
This paper introduces BioNeRF, which combines features extracted from both the camera position and direction into a memory mechanism inspired by cortical circuits [10, 5]. Further, such memory is employed as a context to leverage the information flow into deeper layers. Figure 1 illustrates the BioNeRF architecture, which spans four main modules: (i) Positional Feature Extraction, (ii) Cognitive Filtering, (iii) Memory Updating, and (iv) Contextual Inference.
@@
Cognitive Filtering. This step performs a series of operations from now on called filters that work on the embeddings coming from the previous step [^1] We call those operations filters, for they output a value in the interval $[0,1]$ for each feature coming from the embeddings $\bm{h}_{\Delta}$ and $\bm{h}_{c}$. [^2] Notice such process is inspired in recurrent networks gates’ mechanism, like the LSTM [9].. There are four filters this step derives: (i) density ($\bm{f}_{\Delta}$), color ($\bm{f}_{c}$), memory ($\bm{f}_{\psi}$), and modulation ($\bm{f}_{\mu}$), computed as follows:
$\bm{f}_{\Delta}=\sigma\left(\bm{W}_{\Delta}^{f}{h}_{\Delta}+\bm{b}_{\Delta}^{f}\right),$ (1)
$\bm{f}_{c}=\sigma\left(\bm{W}_{c}^{f}{h}_{c}+\bm{b}_{c}^{f}\right),$ (2)
$\bm{f}_{\psi}=\sigma\left({\bm{W}_{\psi}^{f}\bm{[h}_{\Delta},\bm{h}_{c}]+\bm{b}_{\psi}^{f}}\right),$ (3)
@@
$\bm{f}_{\mu}=\sigma\left({\bm{W}_{\mu}^{f}[\bm{h}_{\Delta},\bm{h}_{c}]+\bm{b}_{\mu}^{f}}\right),$ (4)
where $\bm{W}_{\Delta}^{f}$, $\bm{W}_{c}^{f}$, $\bm{W}_{\psi}^{f}$ and $\bm{W}_{\mu}^{f}$ correspond to the weight matrices for the density, color, memory, and modulation filters, respectively. Additionally, $\bm{b_{\Delta}^{f}}$, $\bm{b_{c}^{f}}$, $\bm{b_{\psi}^{f}}$ and $\bm{b_{\mu}^{f}}$ stand for their respective biases. Moreover, $[\bm{h}_{\Delta},\bm{h}_{c}]$ represents the concatenation of embeddings $\bm{h}_{\Delta}$ and $\bm{h}_{c}$, while $\sigma$ denotes a sigmoid function. The pre-modulation $\bm{\gamma}$ is computed as follows:
@@
Loss Function. Let $r:\Re\times\Re^{3}\rightarrow\Re^{3}$ be a volume rendering technique [11] that computes the pixel color given the volume density and the color. The BioNeRF loss function is defined as follows:
@@
The error is then back-propagated to the model’s previous layers/steps to update its set of weights ($\bm{W}=\{\bm{W}_{\Delta},\bm{W}_{c},\bm{W}_{\Delta}^{f},\bm{W}_{c}^{f},\bm{W}_{\psi}^{f}$,$\bm{W}_{\mu}^{f},\bm{W}_{\gamma},\bm{W}_{\Psi},\bm{W}^{\prime}_{\Delta},\bm{W}^{\prime}_{c}\}$) and biases $\bm{b}=\{\bm{b}_{\Delta},\bm{b}_{c},\bm{b}_{\Delta}^{f},\bm{b}_{c}^{f},\bm{b}_{\psi}^{f}$,$\bm{b}_{\mu}^{f},\bm{b}_{\gamma},\bm{b}_{\Psi},\bm{b}^{\prime}_{\Delta},\bm{b}^{\prime}_{c}\}$.
@@ 4.1 Datasets
We conducted experiments over two well-known datasets concerning view synthesis, i.e., Blender [16] and LLFF [15]:
4.1.1 Blender
Also known as NeRF Synthetic, the Blender comprises eight object scenes with intricate geometry and realistic non-Lambertian materials. Six of these objects are rendered from viewpoints tested on the upper hemisphere, while the remaining two come from viewpoints sampled on a complete sphere.
4.1.2 Local Light Field Fusion
This paper considers $8$ scenes from the LLFF Real dataset, which comprises $24$ scenes captured from handheld cellphones with $20-30$ images each. The authors used a COLMAP structure from motion [19] implementation to compute the poses.
@@ 4.2 Experimental Setup
In the sequence, the memory is used as a context to the consecutive blocks, being concatenated with the camera 3D coordinates to feed the second neural model $M^{\prime}_{\Delta}$, which comprises two dense layers with $256$ neurons and an output layer with a single neuron to predict the volume density $\Delta$. Similarly, the memory is concatenated with the 2D viewing direction $(\theta,\phi)$ to feed the second model $M^{\prime}_{c}$, which is composed of a dense layer with $128$ neurons and an output layer with three neurons to predict the directional emitted color. The network parameters were selected empirically based on the methodology employed in [16], consisting of the Adam optimizer with a learning rate of $45e-$ 4 during $400k$ updates.
@@
The experiments were conducted on an Ubuntu 18 system running on a 2x Intel® Xeon Bronze 3104 processor with $62$GB of memory and an NVIDIA® Tesla T4 GPU. The code was implemented using Python and the PyTorch framework and is available on GitHub [^3] Ommited due to anonymous review..
@@ 5.1 General Evaluation
Table 1 provides the general quantitative results for view synthesis, comparing BioNeRF against state-of-the-art methods considering both synthetic and real image datasets [^4] Values marked with -- describe results not provided by the authors.. The analysis is conducted based on three image reconstruction metrics: PSNR and SSIM, whose objective is to be maximized, and LPIPS, whose lower values denote better results. Notice that bolded numbers represent the most accurate values, while the colors red, orange, and yellow denote the first, second, and third best results, respectively.
Table 1: General results considering both synthetic and real image datasets.
Blender[16] LLFF (Real)[15]
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SRN [21] (NeurIPS’19) $22.26$ $0.846$ $0.170$ $22.84$ $0.668$ $0.378$
LLFF [15] (ACM ToG 2019) $24.88$ $0.911$ $0.114$ $24.13$ $0.798$ $0.212$
NeRF [16] (ACM 21) $31.01$ $0.947$ $0.081$ $26.50$ $0.811$ $0.250$
IBRNet [23] (CVPR’21) $28.14$ $0.942$ $0.072$ $26.73$ $0.851$ $0.175$
Mip-NeRF [3] (ICCV’21) $33.09$ $0.961$ $0.043$- -- --
Mip-NeRF 360 [4] (CVPR’22) $30.34$ $0.951$ $0.060$ $26.59$ $0.846$ $0.168$
TensoRF [6] (ECCV’22) $33.14$ $0.963$ $0.027$ $26,73$ $0.839$ $0.124$
Plenoxels [8] (CVPR’22) $31.71$ $0.958$ $0.049$ $26,29$ $0.839$ $0.210$
Point-NeRF [24] (CVPR’22) $33.31$ $0.978$ $0.027$- -- --
Ref-NeRF [22] (CVPR’22) $\bm{33.99}$ $0.966$ $0.038$- -- --
L2G-NeRF [7] (CVPR’23) $28.62$ $0.930$ $0.070$ $24.54$ $0.750$ $0.200$
Tetra-NeRF [14] (ICCV’23) $32.53$ $\bm{0.982}$ $0.041$- -- --
SpikingNeRF [25] (arXiv’23) $32.45$ $0.956$- -- -- --
Localised-NeRF: [20] (CVPR’24) $33.25$ $0.969$- -- -- --
@@
The results confirm the effectiveness of BioNeRF since it could obtain the best outcomes for all three metrics considering the LLFF dataset, which comprises $8$ scenes from a real environment. BioNeRF also obtained the lowest LPIPS value over the synthetic dataset, outperforming all state-of-the-art results in this context. Such results are highly positive since LPIPS correspond to human perceptual judgments far better than the commonly used metrics like PSNR and SSIM [26].
@@
5.2 Quantitative evaluation
Table 2 presents results concerning the Blender dataset. Regarding the PSNR measure, one can observe that Ref-NeRF obtained the most accurate average results, achieving higher PSNR values over half of the scenes, while Point-NeRF stood out in two of them, and TensoRF and Tetra-NeRF did well in Lego and Ship, respectively. Concerning the SSIM metric, even though Point-NeRF obtained the best results on six out of eight scenes, Tetra-NeRF achieved the best average accuracy overall due to an excellent performance over the Ship scene.
Table 2: Quantitative results considering each scene from Blender dataset.
PSNR $\uparrow$